DVCS TO DO list, Summer 2010

From Hall A Wiki
Jump to: navigation, search

Back to DVCS

This page lists the projects that need to be worked on by the DVCS collaboration, it is not prioritized. Please consider working on this topics. To do so, under each section title state that you are interested on working on this, the date at which you are starting to work on the project, and the status of your work.

For the installation TO-DO list consult: DVCS TO DO list (technical), summer 2010


Slow controls

Anode current ADCs

IOC up and drivers for VMIC3128 initialized, just need to setup the screens and test the ADCs https://hallaweb.jlab.org/dvcslog//ADC/14

LED control

Here is what needs to be done : https://hallaweb.jlab.org/dvcslog//LED/12

Supporting info are: https://hallaweb.jlab.org/dvcslog//LED/7 https://hallaweb.jlab.org/dvcslog//LED/8 https://hallaweb.jlab.org/dvcslog//LED/9

HV controls (100% complete)

The HV crates can be controlled through 3 different methods:

  • telnet: as of 07-07-10: works but the whole method is clumsy
  • ARCNET: in progress, scripts exist by Chris Slominski. Driver was configure. Unable to start the ARCNet driver on the crate so far.

Here is the starting point documentation: https://hallaweb.jlab.org/dvcslog//HV/6

Hardware limits (100% complete)

Need to set the hardware limit to ~900 V using the potentiometer at the back of the crate ( Done by Carlos )


- Unroll and label the 55m RG58 in the hall (100% complete)

- From Carlos (August 12) The whole rack below the DVCS stand will need to be reorganised to replace BNC patch panels by the RG213 patch panels. The HV patch panel should remain, but there is probably no place for the LV power supply (Charles was suggesting to put it where the HV of the PA used to be). You can start setting up patch pannel on the rack and setting up the RG213 cables. The other end of the RG213 cables (HRS hut) should wait until Jack put detectors back in the hut. Jack said that the easiest will be for him to pick up the end which is closest the the L-HRS (so the end which is close to the R-HRS now should go into the DVCS stand).

- From Carlos (August 12) HV cables are in the Hall (spares are still in the Test Lab), and the HV/slow rack is upstream, between R-HRS and the beamline. You can also start cabling these cables (probably easier than the RG213). Check with Ed where he agrees to put the HV rack (I'm sure he doesn't want cables going through the middle of the Hall, at least right now).

Install Calo in the Hall

- From Carlos (August 12) Some lead blocks should be placed in the back side of the base so that the center of gravity stays inside the cartridges while loading the crystals. The lead blocks should stay there until the base is "pinned down" to its final position (after the survey of calorimeter blocks, which will happen probably after I come back). A. Gavalya could tell you more on this, but a couple of lead blocks should be enough. Charles (August 12, later on) Ed thinks we should be ready to stack the crystals next week, probably Tuesday August 17. He will supervise installing any counterweight lead blocks.

- From Carlos (August 12) The few broken PbF2 blocks are now repaired and all blocks are in the cabinet in the Hall. The stacking order is in page of 4 of https://hallaweb.jlab.org/wiki/images/6/6a/Jun28.pdf. Only one change with respect to that picture (trigger number 31 should be block 63 instead of block 23). Shims are already sticked to the upper side of the corresponding blocks, and G10 pieces are also sticked to the corresponding sides of the top row and last column.


07-07-10:Carlos is the contact point with the survey group
[Here] is the description of our survey requirements.

Charles on August 12: The survey crew had just finished (they were unexpectedly available). I don't have final numbers, but the alignment looks good. Height is off by 3 mm, Pitch and yaw require 100-200 micron adjustments. Ed thinks we should be ready to stack the crystals next week, probably Tuesday. He will supervise installing any counterweight lead blocks.

Blue lights

- The blue light system has been delivered in the hall (August 17).
- From Carlos (August 12): The blue light set will be delivered to the machine shop (expected date is Aug 18, next Wed). Alex will be notified, and someone should pick it up and take it to the Hall. The fans are in Chris' office in the VARC, but we don't need them right now (Jack will install them in September).

Implement LUMI read out

Kalyan Allada is working on implementing scaler readout of the LUMI. The lumi will be plugged in the right HRS scalers.



Made them work those little rascals

Triggering system

Charles has been working on this, Magali has started coding (August 17)

Helicity signal and HRS DAQ

Work with Bob Michaels to implement the reading of the helicity in the regular HRS ( Done by Bob, need to implement redundancy checks )

Run start, run stop scripts

Julie starts to work on that as of August 17
Define what we need there and implement

Time Calo versus HRS

Need the calo to be installed and read out in the Hall. During DVCS1, this was done with a paddle, Bodgan W. has lent us a paddle to do this. Bob Michael also suggested to use a pulser to do this.

Setup EPICS events

Julie starts to work on that as of August 17
Do not forget to register the calo anode currents.



  • Instal analysis code on adaq machines done by Julie. The code was installed under adaq:dvcs
  • Setup the online GUI. One should pull out the one that was used by the previous experiment from the MSS, 07-07-10: Sophia will do that


Need to define, create/revive and/or test software scripts needed for offline analysis. The ones that have been identified so far are:

  • Elastic calibration 07-07-10: Asenath is working on that
  • Wave form analysis 07-07-10:Carlos will work on that Done by Carlos
  • Gain monitoring using π0 calibration 07-07-10:Eric has been working on that and will conclude on that shortly


Commissionning run plan

07-07-10: Charles will work on that

Things we discussed briefly so far: do we need to learn how to do the sieve slit analysis before the run starts?

Run plan

07-07-2010: Charles will work on that
Here is the[long term run plan] as of May 2010.
Here is a [first draft] of a more detailed day to day run plan.

How to

Create an HOW-TO: 07-07-10: Eric will work on that

  • based on what already exist from the most recent Hall A experiment (eg [APEX page] )
  • add our own info about how to deal with the calo. [Here] is what was done for the previous DVCS experiment.
  • add info about how to operate the γ detector for the Compton

Shift schedule (100% complete)

The shift schedule is online: [here]

  • Create a shift schedule based on what [PREX used] recently and our existing data base. This is basically ready to go, we are waiting for some more info from JLab about running through the Thanks Giving week-end and through Xmas break.
  • Contact potential run coordinators: 8/8 RC have been identified. Again we are waiting on JLab about running through the Thanks Giving week-end and through Xmas break. Done by Julie.

Run sheet, shift check list

7-19-2010: Eric will do that. Please also look at the list we had for DVCS-1.


07-07-10:Alex will work on that

I gave Pavel our latest run plan.

I have file for COO and ESAD from Doug just need to edit them for us ( will be done by next week )

Still have not found the TOSP. Might start from scratch.